# New Online Communities: Graph Deep Learning on Anonymous Voting Networks to Identify Sybils in Polycentric Governance

### 27 Dec 2023 v3

@@ 6.2   Sybil Resistance and Defence in Depth for Community Design
These efforts, largely emerging from computer science, parallel the digital evolution of Ostrom-style commons management practices emerging in DAOs. DAOs continue to experiment with alternative identity systems, including Proof of Participation (PoP), Decentralized IDs (DIDs), and Self Sovereign IDs (SSIDs). New identity systems that guarantee digital personhood are promising but require complex, unproven infrastructure. This remains an active area of research and development.
@@
The integration of blockchain technology in digital commons has brought about a paradigm shift in how commons security is addressed. Unlike traditional commons, which primarily focus on monitoring and surveillance to prevent illegal resource extraction, digital commons, particularly those managed by DAOs, face a different set of challenges. In response to these evolving security needs, our method does not necessitate structural changes to existing governance systems and requires no additional identity system. Instead, it emphasizes flexible, community-specified security responses to effectively counteract sybil attacks.
@@
If sybils can be accurately and precisely identified, communities can collectively generate rules to deal with their behaviour. As such, effective sybil identification opens up a policy and design space. Ostrom, for instance, recommended leniency against community members who infringe the rules in traditional commons settings. Likewise, digital commons can be flexibly managed by identifying sybils and systematically modulating voting behaviours (by attaching negative karma scores, temporarily blocking spam voting, or even using a form of social naming and shaming). Importantly, our approach is permissive to current and de facto token voting schemes and could—if perfected—permit anonymous token voting while protecting against plutocratic outcomes.
@@ 6.2.1 Defining Sybils and Deanonymizing Social Networks
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is fully supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. Their method fits their goal: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms. Narayanan and Shmatikov’s more recent work on social network deanonymization [35] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [39]) or graph deep learning (Zhou et al. [58]), but these studies still have significant limitations. Paquet-Clouston et al. [39] attempts to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search. Finally, Zhou et al. [58] distinguish between address clustering and account identification, the former we call ‘sybil identification’ and the latter ‘deanonymization.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different and somewhat easier context, is similar to the approach detailed below. The biggest difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search. The second step is necessary to detect sybils because sybils are by definition unknown and with limited connectivity to other voter nodes. Thus we do not have labels for sybils and we cannot a priori determine the number classes for self-supervised learning. A straightforward classification task will not be able to identify sybils.

### 10 Jan 2024 v4

@@ 6.2   Sybil Resistance and Defence in Depth for Community Design
These efforts, largely emerging from computer science, parallel the evolution of Ostrom-style commons management practices emerging in DAOs. DAOs continue to experiment with alternative identity systems, including Proof of Participation (PoP), Decentralized IDs (DIDs), and Self Sovereign IDs (SSIDs). New identity systems that guarantee digital personhood are promising but require complex, unproven infrastructure and mass adoption. This remains an active area of research and development.
Nonetheless, until these new identity systems are perfected and widely adopted, governance systems must still defend against sybil attacks. One such system, already in use at Gitcoin, uses group membership signals to modulate funding payouts for Quadratic Voting (QV), which is a promising method for public funding of digital infrastructure. However, since QV awards matching funds, it is susceptible to coordinated sybil attacks where a user may create multiple identities to extract more funds than appropriate. Specifically, if individuals 1 through n contribute $c_{1}$ through $c_{n}$ to a project, then:
$QV=\left(\sum_{i=1}^{n}\sqrt{c_{i}}\right)^{2}$
Clearly, QV is highly susceptible to sybil attacks because the amplification of identities produces a squared payout.
In a recent work extending the QV system, Miller, Weyl, and Erichsen [34] propose an augmented form of QV that attempts to eliminate the influence of sybils. However, the proposal is not strongly privacy preserving, since it reveals “the intersections of social circles” by leveraging group participation. The authors note that “practically capturing complex social information is non-trivial” and that the Gitcoin solution based on this work (called a “Passport”) captures group membership information by constructing a DID record attached to the user’s wallet. Users are then encouraged to submit credentials, which are added to the DID record, which is itself privacy preserving.
Augmented QV like Gitcoin’s “Passport” appears to offer a unique solution to a particular problem, but the proposed mechanisms are only robust against sybils to the extent that they require attestation to group membership. Miller, Weyl, and Erichsen [34] attempt to find (and formally define) appropriate ways of discounting collusion, based on identifiable group membership, and they offer several sketches for potential defence mechanisms. Three different clustering methods, a pooled centrality measurement, and an ‘eigen match’ are proposed. However, as mentioned, the clustering methods are only effective to the extent that privacy can be reduced by exposing group membership (in theory, a user could construct a Gitcoin ‘Passport’ using only privacy-preserving alternative identity systems, bootstrapping a DID, but this requires social behaviour and mass adoption). The proposed ‘offset match’ is recognised as somewhat crude, since it simply pools users and discounts contributions “to scale correlation scores so that everyone is ‘cared about’ equally strongly.” Finally, the ‘eigen match’ extends the pooling mechanism by focusing on individual behaviours using an adjacency matrix to construct a set of eigenvectors (somewhat similar to calculating Principle Components). However, as I discuss below, due to the anonymous nature of blockchain systems, solutions that rely on edge connections are not able to address singletons, which we assume are common and prevalent. The deep learning method we propose could replace an ‘eigen match’ component and provide more sophisticated QV sybil defence.
@@
Despite these challenges, the integration of blockchain technology and decentralized governance has brought about a paradigm shift in how commons security could be addressed. Unlike traditional commons, which primarily focus on monitoring and surveillance to prevent illegal resource extraction, digital commons, particularly those managed by DAOs, face a different set of challenges. In response to these evolving security needs, our method does not necessitate structural changes to existing governance systems and requires no new identity system. Instead, it emphasizes flexible, community-specified security responses to effectively counteract sybil attacks.
@@
If sybils can be accurately and precisely identified, communities can collectively generate rules to deal with their behaviour. As such, effective sybil identification opens up a policy and design space (including augmenting QV). Ostrom, for instance, recommended leniency against community members who infringe the rules in traditional commons settings. Likewise, digital commons can be flexibly managed by identifying sybils and systematically modulating voting behaviours (by attaching negative karma scores, temporarily blocking spam voting, or even using a form of social naming and shaming). Importantly, our approach is permissive to current and de facto token voting schemes and could—if perfected—permit anonymous token voting while protecting against plutocratic outcomes.
@@ 6.2.1 Defining Sybils and Deanonymizing Social Networks
Prior research, like Backstrom, Dwork, and Kleinberg [4] tackled social network deanonymization, but (like Sweeney) used a dataset that was anonymized after-the-fact. The advantage of these approaches is that the method is fully supervised and has a clear objective. Unfortunately, such methods are not practical against real datasets that achieve their anonymity by cryptographically deriving new addresses. But their methods fit their goals: they propose two graph-based techniques for deanonymization, an active one that requires the prior construction of a subgraph of attacking nodes (prior collusion) to reveal information about neighbours (by walking the graph or making strategic cuts), and a passive method that relies on graph isomorphisms. Narayanan and Shmatikov’s more recent work on social network deanonymization [36] draws inspiration from Backstrom et al., but used graph isomorphisms with auxiliary graphs to link anonymous identities with public identities, a more information rich context and a method that requires a small number of ‘seed’ nodes present in both graphs.
@@
The most effective research on deanonymizing Bitcoin transactions uses real labelled datasets and intuitive machine learning algorithms (Akcora et al. [1] and Paquet-Clouston et al. [40]) or graph deep learning (Zhou et al. [59]), but these studies still have significant limitations. Paquet-Clouston et al. [40] attempt to trace ransomware payments by constructing graphs from 7222 labelled seeds originating from a convenience sample of known ransomware addresses. They construct an address graph and a cluster graph, the latter uses the logic of sybils to combine nodes that have shared inputs (multiple-input clustering heuristics). However, their analysis is purely descriptive and does not attempt to make predictions. Akcora et al. [1] also investigate ransomware transactions, use convenience sample of three datasets (which includes the Paquet-Clouston et al. dataset), and construct a transaction graph and an address graph. They use multiple heuristics, develop graph features, use DBSCAN and k-means algorithms to cluster addresses and a naive similarity search. Finally, Zhou et al. [59] distinguish between address clustering and account identification, the former we call ‘sybil identification’ and the latter ‘deanonymization.’ They construct a trimmed graph of Ethereum transactions, develop node features, and train a graph neural network (GNN) to search subgraph node embeddings using a hierarchical (weighted) graph attention autoencoder. As is typical, they use several pooling layers and non-linear activation layers (LeakyReLu). They use cross-entropy loss for the classifier. They improve the performance of the default GNN by using contrastive subgraph learning. This self-supervised methodology, while in a different and somewhat easier context, is similar to the approach detailed below. The biggest difference is that they classified nodes directly using the GNN, whereas we use develop the embeddings first and then classify using a embeddings search. The second step is necessary to detect sybils because sybils are by definition unknown and with limited connectivity to other voter nodes. Thus we do not have labels for sybils and we cannot a priori determine the number classes for self-supervised learning. A straightforward classification task will not be able to identify sybils.